OpenStudy (anonymous):

A fruit stand has 28 apples, 20 oranges, and 12 pears. What is the ratio of apples to the total pieces of fruit? A. 1 : 4 B. 7 : 15 C. 1 : 2 D. 7 : 8 (I will fan you)

OpenStudy (anonymous):

no of apples / total no of fruits = 28/(28+20+12) = 28/60 ie 28:60 =28/4: 60/4 = 7: 15